Evaluate ∫y0ksin(kx−π)dx where y and k are independent of x, and y is given by the solution to the equation cos(ky)=0.68 .
Answer #1

I think the question is like - de ilk sin fax = ) d2.. Now, sin(kx_19) = - Sim (M-kx) = - Smkz. I = (ksin(kx_1) dx . het = -